OpenStudy (dsfdsf34234):

ONE-THIRD THE PRODUCT OF TWO NUMBERS IS 24. ONE-HALF THE PRODUCT OF THESE SAME TWO NUMBER IS

imqwerty (imqwerty):

let the numbers be \(x\) and \(y\) so \(\Large \frac{x \times y}{3}=24\) so \(\Large \frac{x \times y}{1}=24 \times 3\) so \(\Large \frac{x \times y}{2}=\frac{24 \times 3}{2}\) \(\Large \frac{x \times y}{2}=36\)